OpenStudy (anonymous):

A rectangle has perimeter 46ft and width 8ft. Find the length of the diagonal

OpenStudy (anonymous):

from perimeter find the length. then using length and width u have a right angled traingle whose hypotenuse is the diagonal Using Pythagors theorem find the hypotenuse i.e. the diagonal Do it yourself and i will confirm the answer...... Do you want to try???

OpenStudy (anonymous):

8^2+14^2=c^2

OpenStudy (anonymous):

A rectangle perimeter has 46 ft and width 8 ft. 2L + 2(8) = 46 simplify, divide by 2 L + 8 = 23 L = 23 - 8 L = 15 ft is the length : Find the length of the diagonal (c) c^2 = 15^2 + 8^2 c^2 = 225 + 64 c = c = 17 ft is the diagonal

OpenStudy (anonymous):

you dont use that for only for triangles

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

your welcome :) XD